RC meeting and Run Plan, Thursday Sept. 7, 2023
Return to the RC Meeting Calendar
RC Meeting Information
Time:at 3:30pm today
- Location: Counting house conference room
- ZoomGov Meeting Connection Information
Readiness for Beam
- CG: beam restoration to hall A now projected to start swing Tue Sept 12. We will need to have a shift crew starting on Wed Sept 13th.
- SSG still has a number of items for the HCO but they will probably get to them by the end of the week (after completing Hall C work). Bill worked with them to complete the target FSD work today.
- Ed Folts completing paperwork with everything ready by tomorrow (since it's just name changes no need to resign)
Hall readiness and lockup schedule (Zak Remele)
- ZR: SBS overtemp issue resolved for now; potential for acid flush on Monday is SBS powersupply issue persisits
- ZR: plan to do checklist on Tue (everyone out of the hall by lunch)
Beamline readiness: beamline quads and harps, bcms, bpms, ion chambers (Ciprian Gal)
- Daniel Moser: Beam plans for reviving the target protection systems (the position locks and ramp rate) in place (Bill will review).
- Beam current limit: 45uA.
For each subsystem we ask that the responsible person go over the following points:
- The overall status of your subsystem
- Will your subsystem be ready well before Hall lock-up on Friday or at the latest by Monday morning?
- Does your system being ready depend on your team only or are outside resources required?
- HCal (Jiwan Poudel)
- repaired 3 ch; 2 more chan to go (today and tomorrow) => in the logbook
- Bigbite (Arun Tadepalli)
- already took some cosmic data and things look ok
- Target cell status(Arun/Gordon)
- Gradient field probe measurements complete; magnet behaved
- tomorrow morning Bogdan wants to do the compass measurements and install as much as possible (polarize it over the weekend and do alignment on Monday and dress up the cell)
- Target laser status (alignment, polarization, power etc)
- last step remains to use the alignment laser to get the ladder set up
- Counting House Alarms (Bill Henry)
- Bill tested the alarms and instructions for shift crew response in the wiki (needs further revision?)
- DAQ (Alex Camsonne)
- fixed the SBS crate
- Paul K needs to check the helicity (parity DAQ is fine), AC checked the raster
- plan to run parity DAQ and do online (active) charge feedback for A and C
- warning message on the helicity scaler on the LHRS (may need to be replace the scaler board) => PK and AC will look into it
- GEMs (Holly Szumila-Vance)
- Vimukthi: fixed issue with SBS and BB GEMs (LEMO cables was the culprit - logbook still needs to be made)
- Target magnets and measurements (Bogdan W. and Bill Henry)
- WH: done.
- Moller polarimeter (Don Jones)
- Target motion not working (Ellen looking into it)
- we want to do at least one measurement at the beginning; all systems that could be checked now (besides the target) work as expected
- magnet not cold; expect to get it there end of week (over the weekend)
- SBS magnet and Bigbite magnets ramp and hold?
- seems to work for now
- Corrector magnets:
- need to be tested by Ellen; CG provided max current
- WH: the HV GUI does not populate the EPICS variables; the instructions to restart it don't work;
- GP: the power on the crate was turned on today; will try to power them on from the GUI; and perhaps take a cosmic run
Preliminary run plan
- First, bring the beam safely into the hall (Beam recovery procedure - Jiwan and William Henry https://sbs.jlab.org/wiki/index.php/HOW_TOs#Beam_Recovery_Procedure)
Q: At what stage is the SBS turned ON? WH: will post links to the corrector checkout
- BigBite/single foil location reconstruction (~20 mins?) - BigBite + HCAL detector experts on call consulting analysis experts like Andrew/Sean (magnet on)
- Target performance assessment by target group - Arun (Why hasn’t the target reached xx% polarization?), tell us when we will get there (Target group - Arun) and plan for those hours in between
- Take production data!
List of items in case target isn’t ready:
- Take data on reference cell at 10 atm H2 (portions that add up to 1 day)
- BigBite Sieve slit
- with 8 carbon foils and take optics for 1 hr at 40 uA (Sean/Andrew - confirm time)
- 1 hour for single foil (magnets off) => Chance BBcal HV setting to appropriate values
- Moller measurement depending on availability of people and results from Hall C
In person: Ciprian G, Sanghwa P, Alexandre C, Don J, Jimmy Caylor, Chandan G, Jiwan, Zak
Online: Paul K, Vimukthi H, Brayn M, Jack J, Sean J, Andrew P, Gary P, Todd A, Bhasitha D, Bogdan W,